OpenStudy (mikezack123):

solve using elimation or substitution, which ever. x + 2y = 8 3x + 6y = 12

OpenStudy (kkutie7):

\[x+2y=8\] \[2y=8-x\] \[y=4-\frac{1}{2}x\]

OpenStudy (kkutie7):

\[3x+6(4-\frac{1}{2}x)=12\] you need to solve for x here and plug it into the first equation to find y

OpenStudy (tkhunny):

Not a good plan. x + 2y = 8 3x + 6y = 12 Multiply the top by 3 3x + 6y = 24 3x + 6y = 12 This system is going nowhere. No solution.

OpenStudy (skullpatrol):

The solution set is $$\emptyset$$

OpenStudy (mikezack123):

thank you all

OpenStudy (skullpatrol):

Thanks for asking :-)